Assume that when adults with smart phones are randomly selected, 43% use them on meetings or classes. If 6 adults are randomly selected, find the probability that at least 3 use their phones in class or meetings.


1
Expert's answer
2021-06-21T15:31:20-0400

This is a binomial distribution with n=6, p=0.43.

"P(X\ge3)=1-P(X=0)-P(X=1)-P(X=2)="

"=1-C_6^00.43^0(1-0.43)^6-C_6^10.43^1(1-0.43)^5-C_6^20.43^2(1-0.43)^4="

"=0.5177."